OpenStudy (anonymous):

Asaph took out a loan for $17,550. To pay it back, he will make 60 monthly payments of $437. How much will Asaph pay in interest?

OpenStudy (anonymous):

8670, 60 * 437 = 26220, 26220 - 17550 = 8670

OpenStudy (anonymous):

interest = rate x time ?

OpenStudy (***[isuru]***):

the total amount that Asaph pays = 437 *60 = 26220 the amount that Asaph took as loan = 17 550 the amount that Asap has pay as the interest = 26220 - 17550 = 8670 got it :) ?

OpenStudy (anonymous):

Oh ok! makes sense